Find the scalar and vector product of two vectors A = (3i – 4j + 5k) and B = (-2i + j +3k)?

Grade:11

2 Answers

arbaj adib rahman
8 years ago
Scalar product =(3-4+5)(-2+1+3)=(-6+3+9)+(8-4-12)+(-10+5+15)=6-8+10=8 vector product = -6-4+15=5
Yash raj
5 years ago
a.b=I((4×3)-(5×1))-j((3×-3)-(5×-2))+k((3×1)-(4×-2))
=I(-12-5)-j(-9+10)+k(3+8)
after solving it we get.
=-17i-j+11k